OpenStudy (anonymous):

Multiply: (2x + 4)(3x^2 - 2x - 10)

OpenStudy (chihiroasleaf):

\[\left( 2x+4 \right)\left( 3x ^{2} -2x-10\right) = 2x \left( 3x ^{2} -2x-10 \right)+4\left( 3x ^{2} -2x-10 \right)\] now, use distributive property and simplify

OpenStudy (anonymous):

So the answer should be: 6x^2 - 4x - 20x + 12x^2 - 8x - 40 which breaks down into: 18x^2 - 32x - 40?

OpenStudy (anonymous):

Should have some x^3 in there . . .

OpenStudy (anonymous):

e.g. 2x*3x^2 = 6x^3.

OpenStudy (anonymous):

Oh duh. Okay so. 6x^3 - 4x^2 - 20x + 12x^2 -8x - 40 so breaking down would be 6x^3 + 8x^2 - 28x - 40?

OpenStudy (anonymous):

That's it!

OpenStudy (chihiroasleaf):

Yes.., it's correct... :)

OpenStudy (anonymous):

Thanks for the help, this test seems impossible. <.>